The Excel worksheet form that appears below is to be used to recreate part of the example relating to Turbo Crafters that appears earlier in the chapter. 1. Award: 10.00 points 2. Change all of the numbers in the data area of your worksheet so that it looks like this: 1 Chapter 3: Applying Excel 2 3 Data 4 Allocation base 5 Estimated manufacturing overhead cost 6 Estimated total amount of the allocation base 7 Actual manufacturing overhead cost Machine-hours $ 791,200 92,000 machine-hours $ 773,700 8 Actuil tolal a mount of the locationbase 93,000 machine-hours If your formulas are correct, you should get the correct answers to the following questions. (a) What is the Predetermined overhead rate? (Round your answer to 2 decimal places.) mined overhead per machine hour (b) By how much is the manufacturing overhead underapplied or overapplied? manufacturing overheacd3. Either print or make a copy of your worksheet in your workbook before proceeding. You will need to refer back to this worksheet. Change the estimated total amount of the allocation base to 86,000 machine-hours, but keep everything the same as in Requirement 2. The data area of your worksheet should now look like this: 1Chapter 3: Applying Excel 3 Data 4 Allocation base 5 Estimated manufacturing overhead cost 6 Estimated total amount of the allocation base 7 Actual manufacturing overhead cost 8 Actual total amount of the allocation base chine-hours $791,200 86,000 machine-hours $ 773,700 93,000 machine-hours (a) By how much is the manufacturing overhead underapplied or overapplied?(b) Why is the amount of underapplied (overapplied) manufacturing overhead different from Requirement 2 above? (You may select more than one answer. Single click the box with the question mark to produce a check mark for a correct answers and double click the box with the question mark to empty the box for a wrong answers. Any boxes left with a question mark will be automatically graded as The change in the estimated total amount of the allocation base affected the predetermined overhead rate. The change in the estimated total amount of the allocation base affected the amount of manufacturing overhead applied. Actual manufacturing overhead cost differs from period to period. The estimated manufacturing overhead cost did not change in proportion to the change in the estimated total amount of the allocation base. 4. Change the estimated total amount of the allocation base back to 92,000 machine-hours, so that the data area of you worksheet looks exactly the same as in Requirement 2. Now change the actual manufacturing overhead cost from $773,700 to $778,300. The data area of your worksheet should now look like this: 1Chapter 3: Applying Excel 3 Data 4 Allocation base S Estimated manufacturing overhead cost 6 Estimated total amount of the allocation base 7 Actual manufacturing overhead cost 8 Actual total amount of the allocation base Machine-hours S 791,200 92,000 machine-hours S 778,300 93,000 machine-hours (a) By how much is the manufacturing overhead underapplied or overapplied?(b) Why is the amount of underapplied (overapplied) manufacturing overhead different from Requirement 2 above? (You may select more than one answer. Single click the box with the question mark to produce a check mark for a correct answers and double click the box with the question mark to empty the box for a wrong answers. Any boxes left with a question mark will be automatically graded as incorrect.) The predetermined overhead rate changed. 囝The change in the actual manufacturing overhead cost directly resulted in a comparable change in the amount of underappled overapplied manufacturing overhead. The change in the estimated total amount of the allocation base affected the amount of manufacturing overhead applied
Answer #1

2) a) Predetermined Overhead Rate= estimated manufacturing overhead/ machine hours

= 791200/ 92000= $8.6 per machine hour

b) Manufacturing overheads applied= predetermined rate* actual hours

= 8.6* 93000= $799800

Under/ Over application= Manufacturing overheads recovered- actual overheads

= 799800- 773700= $26100 (over applied)

3) a) Predetermined overhead rate= 791200/ 86000= $9.2 per machine hour

Manufacturing overheads applied= 9.2* 93000= $855600

Under/Over application = 855600- 773700= $81900 ( over applied)

b)The change in the estimated total amount of the allocation base affected the predetermined overhead rate

   The change in the estimated total amount of allocation base affected the amount of manufacturing overhead applied

The estimated manufacturing overhead cost did not change in proportion to the change on the estimated total amount of the allocation base.

4) a) Predetermined overhead rate= $8.6 ( same as requirement 2)

Manufacturing overhead applied= 8.6* 93000= $799800

Under/ Over application= 799800- 778300= $21500(over applied)

b) The change in the actual manufacturing overhead cost directly resulted in a comparable change in the amount of over-applied overhead